# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ 2.3. Deterministic error analysis
Let $\gamma$ be an $\alpha$-power spectral density. Then, there is a constant $C>0$ such that for all $f\in L^{2}(\mathcal{M})$
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $\|_{L^{2}(\mathcal{M})}\leq\begin{cases}Ch^{2}\|f\|_{L^{2}(\mathcal{M})},&\alpha>1,\\ Ch^{2}|\log h|\|f\|_{L^{2}(\mathcal{M})},&\alpha=1,\\ Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})},&\alpha<1.\end{cases}$
To prove this proposition we rely on a representation of functions of operators based on Cauchy–Stieltjes integrals, which are constructed as follows. Since the sesquilinear form defined by $\mathsf{A}_{\mathcal{M}}$ is continuous and coercive, the associated operators $\mathcal{L}$ and $\mathcal{L}_{h}$ are sectorial with some (common) angle $\theta\in(0,\pi/2)$ (Yagi, 2010, Theorem 2.1). Therefore, the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ are contained in the complement of the set $G_{\theta}=\{z\in\amsmathbb{C},\theta\leq\arg(z)\leq\pi\}$, as illustrated in Figure 3(a), and the following inequalities are satisfied for any $z\in G_{\theta}$ (cf. Yagi (2010, Equation (2.2))):
$\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v\|_{L^{2}(\mathcal{M})},\quad v\in L^{2}(\mathcal{M}),$
@@
where $C_{\theta}>0$ is a generic constant. Note in particular that by definition, $G_{\theta}$ is contained in the resolvent sets of $\mathcal{L}$ and $\mathcal{L}_{h}$, and that any power spectral density $\gamma$ is bounded, holomorphic and satisfies the inequality $|z|^{\alpha}|\gamma(z)|\lesssim 1$ for any $z\in G_{\theta}$. Hence, the operators $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$ can be defined as functional calculi of the operators $\mathcal{L}$ and $\mathcal{L}_{h}$ as (Yagi, 2010, Chapter 16, Section 1.2) by
@@
where $\Gamma\subset\amsmathbb{C}$ is any integral contour surrounding the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ and contained in $G_{\theta}$. In particular, these new definitions of functions of operators are independent of the choice of $\Gamma$, and coincide with the spectral definitions previously introduced in Equation 8 (cf. e.g. (Yagi, 2010, Remark 2.7)).
@@
where $\Gamma_{+}$ is parametrized by $g_{+}(t)=te^{i\theta}$ for $t\in(\infty,\delta_{0}]$, $\Gamma_{0}$ by $g_{0}(t)=\delta_{0}e^{it}$ for $t\in[\theta,-\theta]$ and $\Gamma_{-}$ by $g_{-}(t)=te^{i\theta}$ for $t\in[\delta_{0},\infty)$ with $\delta_{0}<\delta/2$ (see Figure 3(b) for an illustration). We use this contour to prove Proposition 2.3, while relying on the following bounds for the resolvent error along $\Gamma$. The proof is included in Appendix A and is an adaption of results from Fujita and Suzuki (1991) and Bonito et al. (2024).
@@
For any $z\in\Gamma$, $f\in L^{2}(\mathcal{M})$, and $\beta\in[0,1]$,
$\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq Ch^{2\beta}|z|^{-(1-\beta)}.$
@@
Let $f\in L^{2}(\mathcal{M})$. For any $z\in\Gamma$, set $e_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f$, which yields, using the integral representations of $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$,
$\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)e_{h}(z)\,\mathrm{d}z.$
@@
$\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $=\frac{-1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{+}(t))e_{h}(g_{+}(t))g^{\prime}_{+}(t)\,\mathrm{d}t-\frac{1}{2\pi i}\int_{-\theta}^{\theta}\gamma(g_{0}(t))e_{h}(g_{0}(t))g_{0}^{\prime}(t)\,\mathrm{d}t$
$\quad+\frac{1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{-}(t))e_{h}(g_{-}(t))g^{\prime}_{-}(t)\,\mathrm{d}t,$
@@
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
$\quad\leq\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{+}(t))|\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{-\theta}^{\theta}|\gamma(g_{0}(t))|\|e_{h}(g_{0}(t)\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
$\qquad+\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{-}(t))|\left\|e_{h}(g_{-}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
@@
Let us start by bounding $I_{0}$. We apply Equation 7 and Lemma 2.4 with $\beta=1$ to obtain
@@
$I_{0}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}|g_{0}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}\delta_{0}^{-\alpha}\,\mathrm{d}t\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}.$
@@
To bound $I_{+}$, we distinguish between the three cases $\alpha>1$, $\alpha<1$ and $\alpha=1$. When $\alpha>1$, we apply Equation 7 and Lemma 2.4 with $\beta=1$ to get
$I_{+}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|g_{+}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\int_{\delta_{0}}^{\infty}|t|^{-\alpha}\,\mathrm{d}t=C\frac{\delta_{0}^{-(\alpha-1)}}{\alpha-1}h^{2}\|f\|_{L^{2}(\mathcal{M})}.$
For $\alpha<1$, recall that $h<h_{0}<1$ and assume without loss of generality that $h_{0}^{-2}>\delta_{0}$. We split $I_{+}=I_{+1}+I_{+2}$ with
@@
$I_{+1}=\int_{\delta_{0}}^{h^{-2}}|\gamma(g_{+}(t))|\,\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,I_{+2}=\int_{h^{-2}}^{\infty}|\gamma(g_{+}(t))|\,\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t.$
Using Equation 7 and Lemma 2.4 with $\beta=1$ we bound $I_{+1}$ by
$I_{+1}$ $\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}|g_{+}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}|t|^{-\alpha}\,\mathrm{d}t$
$\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{0}^{h^{-2}}|t|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\frac{h^{-2(1-\alpha)}}{1-\alpha}\leq Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})}.$
We proceed in the same manner to bound $I_{+2}$, but use Lemma 2.4 with $\beta=0$,
@@
$I_{+2}\leq C\|f\|_{L^{2}(\mathcal{M})}\int_{h^{-2}}^{\infty}|g_{+}(t)|^{-\alpha}|g_{+}(t)|^{-1}\,\mathrm{d}t=C\|f\|_{L^{2}(\mathcal{M})}\int_{h^{-2}}^{\infty}|t|^{-\alpha-1}\,\mathrm{d}t\leq Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})}.$
For $\alpha=1$ and the same splitting $I_{+}=I_{+1}+I_{+2}$ we obtain for $I_{+1}$ with Lemma 2.4 for $\beta=1$
$I_{+1}$ $\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}t^{-1}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(\log(h^{-2})-\log\delta_{0})$
$\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(|\log(h^{-2})|+|\log\delta_{0}|)\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(|\log h|+|\log\delta_{0}|).$
Assuming again $h<h_{0}$ and $h_{0}^{-2}>\delta_{0}$, we finally bound
$I_{+1}$ $\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}\left(1+\left|\frac{\log\delta_{0}}{\log h}\right|\right)\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}\left(1+\left|\frac{\log\delta_{0}}{\log h_{0}}\right|\right)$
$\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}.$
@@
The bound for $I_{+2}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}$ is obtained in the same way as in the case $\alpha>1$. Hence, $I_{+}\leq Ch^{2}|\log h|\leavevmode\nobreak\ \|f\|_{L^{2}(\mathcal{M})}$ for $\alpha=1$.
Finally, note that by symmetry, $I_{-}$ satisfies the same bounds as $I_{+}$, which finishes the proof adding up all terms. ∎
@@
Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. There exists a constant $C>0$ such that, for any $\tilde{f}\in S_{h}^{\ell}$,
(24) $\left\|\left(\gamma(\mathcal{L}_{h})\tilde{f}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\tilde{f}^{-\ell})\right\|_{L^{2}(\mathcal{M}_{h})}\leq Ch^{2}\,\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})}.$
This proposition can be seen as an extension of Bonito et al. (2024, Lemma 4.4) relying on extensions of the estimates proven in Bonito et al. (2024, Lemma A.1). Its proof is similar and therefore postponed to Appendix B.

### v2

@@ 2.3. Deterministic error analysis
From now on, let us make the following assumption on the mesh size $h$. Let $\delta_{0}\in(0,\delta/2)$ be fixed and arbitrary and let $h_{0}\in(0,1)$ such that $h_{0}^{-2}>\delta_{0}$ and $|\log h_{0}|>1$. We now assume for the remainder of the paper n that mesh size satisfies $h\in(0,h_{0})$.
@@
Let $\gamma$ be an $\alpha$-amplitude spectral density. There exists some constant such that for any $h\in(0,h_{0})$, for all $f\in L^{2}(\mathcal{M})$ and for any $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$, we have
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $\|_{L^{2}(\mathcal{M})}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
where $C_{\alpha+p}(h)=|\log h|$ if $\alpha+p\leq 1$, and $C_{\alpha+p}(h)=1$ otherwise.
To prove this proposition we rely on a representation of functions of operators based on Cauchy–Stieltjes integrals, which are constructed as follows. Since the sesquilinear form defined by $\mathsf{A}_{\mathcal{M}}$ is continuous and coercive, the associated operators $\mathcal{L}$ and $\mathcal{L}_{h}$ are sectorial with some (common) angle $\theta\in(0,\pi/2)$ [28, Theorem 2.1]. Therefore, the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ are contained in the complement of the set $G_{\theta}=\{z\in\amsmathbb{C},\theta\leq\arg(z)\leq\pi\}$, as illustrated in Figure 3(a), and the following inequalities are satisfied for any $z\in G_{\theta}$ (cf. [28, Equation (2.2)]):
(20) $\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v\|_{L^{2}(\mathcal{M})},\quad v\in L^{2}(\mathcal{M}),$
@@
where $C_{\theta}>0$ is a generic constant. Note in particular that by definition, $G_{\theta}$ is contained in the resolvent sets of $\mathcal{L}$ and $\mathcal{L}_{h}$, and that any amplitude spectral density $\gamma$ is bounded, holomorphic and satisfies the inequality $|z|^{\alpha}|\gamma(z)|\lesssim 1$ for any $z\in G_{\theta}$. Hence, the operators $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$ can be defined as functional calculi of the operators $\mathcal{L}$ and $\mathcal{L}_{h}$ as [28, Chapter 16, Section 1.2] by
@@
where $\Gamma\subset\amsmathbb{C}$ is any integral contour surrounding the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ and contained in $G_{\theta}$. In particular, these new definitions of functions of operators are independent of the choice of $\Gamma$, and coincide with the spectral definitions previously introduced in Equation 8 (cf. e.g. [28, Remark 2.7]).
@@
where $\Gamma_{+}$ is parametrized by $g_{+}(t)=te^{i\theta}$ for $t\in(\infty,\delta_{0}]$, $\Gamma_{0}$ by $g_{0}(t)=\delta_{0}e^{it}$ for $t\in[\theta,-\theta]$ and $\Gamma_{-}$ by $g_{-}(t)=te^{i\theta}$ for $t\in[\delta_{0},\infty)$ with $\delta_{0}<\delta/2$ (see Figure 3(b) for an illustration). We use this contour to prove Proposition 2.4, while relying on the following bounds for the resolvent error along $\Gamma$. The proof is included in Section B.3 and is an adaption of results from [13, 6, 7].
@@
There exists some constant such that for any $h\in(0,h_{0})$, for any $z\in\Gamma$, any $f\in L^{2}(\mathcal{M})$, for any $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$, for any $\beta\in[0,1]$ such that $p\in[0,(1+\beta)/2]$,
(24) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-P_{h}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\lesssim h^{2\beta}|z|^{-(1+p-\beta)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
@@
Let $h\in(0,h_{0})$, $f\in L^{2}(\mathcal{M})$ and let $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$. We can therefore introduce $u=\mathcal{L}^{p}f\in L^{2}(\mathcal{M})$. First, note using the triangle inequality, we have
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$ $\leq\|\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
$\quad+\|P_{h}\gamma(\mathcal{L})f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}=S_{1}+S_{2}$
where we take $S_{1}=\|\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$ and $S_{2}=\|P_{h}\gamma(\mathcal{L})f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$. We now bound these two terms.
For the term $S_{2}$, let us introduce the function $\widehat{\gamma}$ defined as $\widehat{\gamma}(\lambda)=\gamma(\lambda)\lambda^{-p}$. In particular, note that since $\gamma$ is an $\alpha$-amplitude spectral density, $\widehat{\gamma}$ is an $(\alpha+p)$-amplitude spectral density, and we have, by definition of $u$,
$S_{2}=\|(I-P_{h})\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}=\|(I-P_{h})\widehat{\gamma}(\mathcal{L})u\|_{L^{2}(\mathcal{M})}$
We then use the Bramble–Hilbert lemma (cf. Lemma B.4) with $t=2\min\{\alpha+p;1\}\in(0,2]$, while noting that, following Remark 2.2, $\|\mathcal{L}^{t/2}(\widehat{\gamma}(\mathcal{L})u)\|_{L^{2}(\mathcal{M})}\lesssim\|u\|_{L^{2}(\mathcal{M})}<\infty$ (since $t/2\leq\alpha+p$):
$S_{2}\lesssim h^{t}\|\mathcal{L}^{t/2}(\widehat{\gamma}(\mathcal{L})u)\|_{L^{2}(\mathcal{M})}\lesssim h^{t}\|u\|_{L^{2}(\mathcal{M})}=h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
Let us now bound the term $S_{1}$. For any $z\in\Gamma$, set $\mathcal{F}_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}f-P_{h}(z-\mathcal{L})^{-1}f$, which yields, using the integral representations of $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$,
$\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f$ $=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}_{h}(z)\,\mathrm{d}z.$
@@
$\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f$ $=\frac{-1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{+}(t))\mathcal{F}_{h}(g_{+}(t))g^{\prime}_{+}(t)\,\mathrm{d}t$
$\quad-\frac{1}{2\pi i}\int_{-\theta}^{\theta}\gamma(g_{0}(t))\mathcal{F}_{h}(g_{0}(t))g_{0}^{\prime}(t)\,\mathrm{d}t$
$\quad+\frac{1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{-}(t))\mathcal{F}_{h}(g_{-}(t))g^{\prime}_{-}(t)\,\mathrm{d}t,$
@@
$S_{1}=\|\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
$\quad\leq\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{+}(t))|\left\|\mathcal{F}_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{-\theta}^{\theta}|\gamma(g_{0}(t))|\|\mathcal{F}_{h}(g_{0}(t))\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
$\qquad+\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{-}(t))|\left\|\mathcal{F}_{h}(g_{-}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
@@
Let us start by bounding $I_{0}$. We apply Equation 7 and Equation 24 (with $\beta=1$) to obtain
@@
$I_{0}$ $\lesssim\int_{-\theta}^{\theta}|g_{0}(t)|^{-\alpha}\leavevmode\nobreak\ h^{2}|g_{0}(t)|^{-p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}\delta_{0}^{-(\alpha+p)}\,\mathrm{d}t$
$\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
@@
To bound $I_{+}$, we distinguish between the three cases $\alpha+p>1$, $\alpha+p<1$ and $\alpha+p=1$. When $\alpha+p>1$, we apply Equation 7 and Equation 24 (with $\beta=1$) to get
$I_{+}$ $\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|g_{+}(t)|^{-(\alpha+p)}\,\mathrm{d}t=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|t|^{-(\alpha+p)}\,\mathrm{d}t$
$=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\frac{\delta_{0}^{-(\alpha+p-1)}}{\alpha+p-1}.$
Hence, we can conclude that if $\alpha+p>1$, we have $I_{+}\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$.
For $\alpha+p<1$, recall that $10<h<h_{0}<$ 1 for some fixed $h_{0}$ such that $h_{0}^{-2}>\delta_{0}$. We then introduce $\tilde{\alpha}=(\alpha+p)/\alpha$ and split $I_{+}=I_{+1}+I_{+2}$ with
@@
$I_{+1}=\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|\gamma(g_{+}(t))|\,\left\|\mathcal{F}_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,$
@@
$I_{+2}=\int_{h^{-2\tilde{\alpha}}}^{\infty}|\gamma(g_{+}(t))|\,\left\|\mathcal{F}_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t.$
Let then $e_{h}=(1-(\alpha+p))/|\log h|\in(0,1-(\alpha+p))$ and $\beta=\alpha+p+e_{h}\in(0,1)$. Note in particular that $p\in[0,(1+\beta)/2]$ since $(1+\beta)/2-p=(1-p+\alpha+e_{h})/2>0$. We can therefore use Equation 7 and Equation 24 to bound $I_{+1}$ by
$I_{+1}$ $\lesssim h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|g_{+}(t)|^{-(\alpha+1+p-\beta)}\,\mathrm{d}t=h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|t|^{-(1-e_{h})}\,\mathrm{d}t$
$\lesssim h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{0}^{h^{-2\tilde{\alpha}}}|t|^{-(1-e_{h})}\,\mathrm{d}t=h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\frac{h^{-2\tilde{\alpha}e_{h}}}{e_{h}}$
$=h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\frac{h^{-2(\tilde{\alpha}-1)e_{h}}}{e_{h}}\lesssim h^{2(\alpha+p)}|\log h|\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
since $h^{-2\tilde{\alpha}e_{h}}=\exp(2\tilde{\alpha}(1-(\alpha+p)))\lesssim 1$. We proceed in the same manner to bound $I_{+2}$, using this time Equation 24 (with $\beta=p$):
$I_{+2}$ $\lesssim h^{2p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{h^{-2\tilde{\alpha}}}^{\infty}|g_{+}(t)|^{-(\alpha+1)}\,\mathrm{d}t=h^{2p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{h^{-2\tilde{\alpha}}}^{\infty}|t|^{-(\alpha+1)}\,\mathrm{d}t$
$\lesssim h^{p}h^{2\tilde{\alpha}\alpha}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\lesssim h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
Hence, we conclude that if $\alpha+p<1$, we have $I_{+}\lesssim|\log h|\leavevmode\nobreak\ h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$.
Finally, for $\alpha+p=1$ we repeat the same approach as for the case $\alpha+p<1$. On the one hand we use Equation 24 (with $\beta=1$) to get for the term $I_{+1}$ the bound
$I_{+1}$ $\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|t|^{-1}\,\mathrm{d}t=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}(\log(h^{-2\tilde{\alpha}})-\log\delta_{0})$
$\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}|\log h|,$
where we used the fact that $h$ is upper-bounded by a fixed constant $h_{0}$ satisfying $h_{0}^{-2}>\delta_{0}$ to derive the last inequality. On the other hand, we once again use Equation 24 with $\beta=p$ to bound the term $I_{+2}$
$I_{+2}$ $\lesssim h^{2p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{h^{-2\tilde{\alpha}}}^{\infty}|t|^{-(\alpha+1)}\,\mathrm{d}t\lesssim h^{p}h^{2\tilde{\alpha}\alpha}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$
$\lesssim h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
@@
Hence, we have $I_{+}\lesssim h^{2}|\log h|\leavevmode\nobreak\ \|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$ when $\alpha+p=1$.
Putting together the three cases, we can therefore conclude that
$I_{+}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
where $C_{\alpha+p}(h)=|\log h|$ if $\alpha+p\leq 1$, and $C_{\alpha+p}(h)>1$ if $\alpha+p\neq 1$. Finally, note that by symmetry, $I_{-}$ satisfies the same bounds as $I_{+}$, meaning that we also get $I_{-}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$.
In conclusion, putting together the estimates for $I_{0}$, $I_{+}$ and $I_{-}$ we get
$S_{1}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
and putting together the estimates pf $S_{1}$ and $S_{2}$ yields
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
which concludes the proof. ∎
@@
Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$. Then, there exists some constant such that for any $h\in(0,h_{0})$, for any $\tilde{f}\in S_{h}^{\ell}$,
(26) $\left\|\left(\gamma(\mathcal{L}_{h})\tilde{f}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\tilde{f}^{-\ell})\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}\,\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})}.$
This proposition can be seen as an extension of [6, Lemma 4.4] relying on extensions of the estimates proven in [6, Lemma A.1]. Its proof is similar and therefore postponed to Appendix B.